OpenStudy (dinamix):

i make some challenge , who answer my question less then 5 min without use google , find this integral or original function of this f(x) = 1/(sin(x)

OpenStudy (empty):

Oh find the integral of \(f(x)=\csc(x)\)? This is a fun problem it has a neat trick. :D

OpenStudy (dinamix):

\[\int\limits_{}^{}\frac{ dx}{ \sin (x) } \]

ganeshie8 (ganeshie8):

\[\int \dfrac{1}{\sin x}\, dx = \int \dfrac{\sin x}{1-\cos^2 x}\, dx = -\int \dfrac{1}{1-u^2}\, du =\cdots \]

OpenStudy (empty):

I guess this trick isn't as systematic as ganehie's but this is how I first learned to do this one: \[\int \csc x dx = \int \csc x \frac{\csc x + \cot x}{\csc x + \cot x} dx = \int \frac{\csc ^2 x + \csc x\cot x}{\csc x + \cot x} dx \cdots \]

OpenStudy (dinamix):

@ganeshie8 i know the answer and i have easy method

ganeshie8 (ganeshie8):

I know that you know :)

OpenStudy (dinamix):

lol nice dude

OpenStudy (empty):

Ok can we solve this with our arms tied behind our backs? You have to find a new way to solve this.

ganeshie8 (ganeshie8):

do you have any other ways... actually the previous trick works for \(\csc^3x\) too but you will need to love partial fractions to use it : \[\int \dfrac{1}{\sin^3 x}\, dx = \int \dfrac{\sin x}{(1-\cos^2 x)^2}\, dx = -\int \dfrac{1}{(1-u^2)^2}\, du =\cdots \]

OpenStudy (dinamix):

@ganeshie8 so what your solution

ganeshie8 (ganeshie8):

I like this one : \[\int \dfrac{1}{\sin x}\, dx = \int\dfrac{1+\tan^2(x/2)}{2\tan(x/2)}\,dx = \int \dfrac{1}{u}\,du = \log\left(\tan(x/2)\right)+C \]

OpenStudy (dinamix):

yeah this is !! u are amazing mate

ganeshie8 (ganeshie8):

Ahh thought you have some other clever way..

OpenStudy (dinamix):

yup i have other way

ganeshie8 (ganeshie8):

please do share xD

OpenStudy (dinamix):

its good challenge or no ?

ganeshie8 (ganeshie8):

sure it is! idk about others, but it really challenged me because i keep forgetting the antiderivatives of cscx and secx and rely on wolfram too much

OpenStudy (empty):

Yeah same, I had to check real quick: \[\frac{d}{dx} (\sin x)^{-1} = - \sin^{-2}x \cos x = -\csc x \cot x\]

OpenStudy (empty):

Also I was trying to see if I could solve it using the infinite product form: \[\large \int \csc x dx = \int \frac{1}{x}\prod_{n=1}^\infty \frac{1}{1-\left(\frac{x}{n \pi} \right)} dx\] But nothing really stands out to me.

OpenStudy (dinamix):

but -cscxcotx= log(tan(x/2) + c or no

OpenStudy (dinamix):

http://prntscr.com/88rahv , @ganeshie8 ,@Empty i hope u to understand my solution

OpenStudy (dinamix):

@Empty

OpenStudy (dinamix):

@IrishBoy123

OpenStudy (irishboy123):

just watching but now that you ask!!! i looked at this integral recently and found this article: https://en.wikipedia.org/wiki/Integral_of_the_secant_function i am fascinated by the history of maths. often such history is about how something gets discovered and used before it is really understood, eg calculus, complex numbers in this case, the secant integral was just a big thing [secant but same difference] .... and the solution would at the time have been pure rocket science.....

OpenStudy (irishboy123):

just my 2 cents :p

OpenStudy (dinamix):

@IrishBoy123 did u see my solution

OpenStudy (dinamix):

@dan815

OpenStudy (irishboy123):

yes, it's the same as @ganeshie8 posted above sweet as a nut!

OpenStudy (dinamix):

yup i know this is

OpenStudy (ikram002p):

@IrishBoy123 nuts is something salty xD

OpenStudy (irishboy123):

lol!!!! OK, sweet as a nut that has been dipped in honey for a week!

OpenStudy (dinamix):

@IrishBoy123 xD

OpenStudy (ikram002p):

still my joke is better xD